# From Simulation to Strategy: Automating Personalized Interaction Planning for Conversational Agents

### v1

@@ Abstract
Amid the rapid rise of agentic dialogue models, realistic user-simulator studies are essential for tuning effective conversation strategies. This work investigates a sales-oriented agent that adapts its dialogue based on user profiles spanning age, gender, and occupation. While age and gender influence overall performance, occupation produces the most pronounced differences in conversational intent. Leveraging this insight, we introduce a lightweight, occupation-conditioned strategy that guides the agent to prioritize intents aligned with user preferences, resulting in shorter and more successful dialogues. Our findings highlight the importance of rich simulator profiles and demonstrate how simple persona-informed strategies can enhance the effectiveness of sales-oriented dialogue systems.
@@ I Introduction
To enhance system performance, a deeper understanding of user behavior within these interactions is essential. By analyzing patterns in user responses, preferences, and conversational dynamics, it is possible to tailor dialogue strategies that foster engagement and drive successful outcomes. Several studies have investigated how different user profiles influence conversational dynamics and how to design more engaging interactions [1, 2, 3, 4]. For example, Zhang et al. [1] were among the first to incorporate user profiles into dialogue systems, enabling the generation of more personalized and engaging user responses. In addition, Cho et al.[2] proposed a personalized dialogue generation framework that implicitly detects user persona from dialogue history via conditional variational inference. Their model enables the agent to generate responses that are better aligned with the user’s latent persona, leading to improved engagement and personalization without requiring explicit user profiles. Furthermore, Wang et al. [3] introduced USP, a user simulation framework that models implicit user traits—such as personality, speaking style, and goals—directly from dialogue data. By leveraging these latent profiles, their system can generate more human-like and behaviorally consistent user interactions in multi-turn conversations.
In this study, we further extend this line of research by investigating the importance of user behavior modeling in the context of agentic dialogue systems and exploring methodologies for leveraging these insights to improve system efficacy. Specifically, we focus on a sales-oriented dialogue system, SalesAgent, proposed by Chang and Chen [5], which aims to proactively shift the dialogue topic toward the user’s potential interests and subsequently elicit explicit expressions of interest in the agent’s recommendations. We conduct a thorough analysis of a broad range of user profiles, including gender, age, and occupation, to examine how different user personas impact conversational flow.
@@ II-A Role-playing Research in LLMs
The role-playing capabilities of large language models (LLMs) have attracted growing attention in recent research [6, 7, 8]. Park et al.[6] introduced Generative Agents with dynamic memory and reflective reasoning to simulate human-like interactions, while Shao et al.[8] proposed Character-LLM, which leverages curated profiles and reconstructed experiences for more consistent persona modeling. Beyond general frameworks, domain-specific role-playing has been explored in software engineering [9, 10, 11], where MetaGPT [9], ChatDev [10], and self-collaboration frameworks [11] enable multi-agent collaboration through structured communication and role-based expertise. In the medical domain, Wu et al.[12] introduced Diagnostic-Reasoning Chain-of-Thought (DR-CoT) prompting for iterative doctor-patient interactions, and Tang et al.[13] proposed MedAgents, a multi-agent framework for collaborative medical reasoning.
@@ II-B SalesAgent
SalesAgent [5] is a dialogue agent designed to bridge the gap between open-domain chitchat and task-oriented dialogues in sales-oriented settings. It builds upon the foundation established by SalesBot [14], which introduced the first large-scale dataset of dialogues that naturally transition from social conversation to task-oriented interactions. SalesBot demonstrated that triggering business opportunities often requires agents to engage users in casual conversation and gradually detect implicit intents. However, early versions of SalesBot suffered from abrupt transitions and limited coherence.
To address these limitations, SalesAgent is trained on an improved dataset, SalesBot 2.0 [15], which enhances conversational flow by incorporating longer and smoother transitions and leveraging commonsense knowledge from large language models (LLMs). Inspired by recent advances in prompting paradigms such as chain-of-thought reasoning [16] and ReAct prompting paradigms [17], SalesAgent introduces explicit strategy-guided reasoning into the dialogue process via chain-of-thought (CoT) prompting. At each turn, it generates two outputs: a thought, which articulates its internal reasoning about the conversational state and user intent, and a response, which is the surface utterance produced for the user.
@@
In addition to CoT-driven strategy control, recent studies have highlighted the importance of modeling user diversity in evaluating sales agents. Cheng et al. [18] proposed MBTI-based persona simulators to systematically analyze how different personality traits affect interaction outcomes with SalesAgent. Their experiments showed that user traits such as openness and extraversion influenced both task success rates and conversational dynamics, underscoring the need for adaptable dialogue strategies.
@@ III-A Persona Attribute Definition
Each user simulator is assigned a predefined persona composed of four attributes: gender, age, occupation, and personality [19, 20].
@@
Occupations are selected following the International Standard Industrial Classification of All Economic Activities (ISIC). From the full list of 21 sections, we chose six representative sectors and sampled four occupations from each. The selected sectors are: Agr, Info , Fin , Edu (Education), Heal , and Arts. A detailed list of these sectors and sampled occupations is provided in Appendix C.
Personality traits are modeled using the Myers-Briggs Type Indicator (MBTI) [21], a widely adopted psychological framework that categorizes individuals into 16 types across four dichotomous dimensions. In this work, we consider the eight individual MBTI traits: Extraversion (E), Introversion (I), Sensing (S), Intuition (N), Thinking (T), Feeling (F), Judging (J), and Perceiving (P). These attributes collectively define the persona, influencing the user simulator’s goals, dialogue behavior, and interaction patterns.
@@ III-B Persona Dataset Formation
For each fixed attribute condition, we generate 20 unique user personas. Persona generation is guided through structured prompting of LLaMA-3.1-8B-Instruct [22], enabling coherent and diverse persona instantiation aligned with the specified constraints. The full prompt and an example of a generated user profile is presented in Appendix A.
@@ III-C Simulation
Each simulation involved a dialogue between two models: a user simulator and the SalesAgent. The user simulator was responsible for generating persona-driven user utterances, while the SalesAgent produced both internal reasoning (thought) and outward responses at each turn. The prompt used for user model is fixed across all experiments and is presented in Appendix B.
@@ III-D Metrics
Inspired by Cheng et al. [18], two metrics were adopted to evaluate the performance of the sales agent, allowing assessment not only of the success rate but also the understanding of the interaction efficiency. We also constructed two visualizations to further analyze the user and the sales agent’s behavior across intents.
@@ IV Persona Attribute Analysis
Simulations were conducted for each user attribute—age, gender, and occupation—resulting in a total of 9000 conversations and at most 180,000 dialogue turns. The user model was LLaMA-3.1-8B-Instruct [22], and the agent model was SalesAgent. The resulting metrics are summarized in Table I. In the following subsections, we discuss our findings with respect to age, gender, and occupation.
TABLE I: SalesAgent performance when interacting with different user types. : $p<0.05$; : $p<0.10$ in significance tests.
@@ IV-A Age
To assess the statistical significance of these patterns, we conducted an analysis of variance (ANOVA) across the four age groups. As shown in Table I, the success rate yielded a $p=0.02$, indicating a statistically significant difference ($p<0.05$) in persuasion effectiveness among age groups.
@@ IV-C Occupation
TABLE II: Mapping of strategy prompts for different user occupations.
@@ V Occupation-Based Strategy for SalesAgent
To this end, we designed a lightweight occupation-based strategy in which SalesAgent prioritizes the two intents most frequently associated with each occupation, as identified through prior simulation analysis (Table II). The goal was to evaluate whether this automatically derived strategy meaningfully improves success rates and conversational efficiency when applied to new user interactions, under the assumption that the user’s occupation is known.
To validate the generalization of this approach, we deliberately used different models for simulation and testing. The strategy was derived from intent distributions observed in the first experiment (using one simulator), but in this experiment, user inputs were generated by Qwen3-8B [23], a different LLM, to ensure that the strategy was not overfitted to a particular simulator for generalizability. This design tests whether patterns learned from one simulated population can transfer to interactions with different user models.
In addition to testing strategy effectiveness, we sought to validate the modularity of this approach. We implemented a lightweight pipeline where SalesAgent provides high-level dialogue planning (i.e., thought generation), while a vanilla LLM generates the final user-facing response. Specifically, Qwen3-8B simulates user input, SalesAgent generates intermediate thoughts, and Mistral-7B-Instruct-v0.3 [24] produces the final response conditioned on both the thought and the strategy input. This architecture highlights how strategy augmentation can be flexibly applied to off-the-shelf LLMs. The prompting setup is detailed in Appendix D.
@@
TABLE III: Performance of the occupation-based SalesAgent (w/o vs. w/ strategies).
@@
Additional Metric In this experiment, we introduced another metric, guided continuation ratio, a slight modification of the continuation ratio proposed by Cheng et al. [18]. This metric measures the proportion of the sales agent’s internal thoughts labeled as “continue” that occur immediately after the point at which the agent decides to pivot the conversation toward a specific topic. This metric helps evaluate the aggressiveness of the agent: the higher the ratio, the more aggressive the agent is.
Results and Insights. Table III demonstrates that the occupation-based strategy yielded notable improvements in success rate, with sections Edu, Heal, and Arts showing gains exceeding 40%. Moreover, both the guided-continuation ratio and the average number of turns declined across most occupations, for both successful and unsuccessful dialogues. This suggests that the strategy enables the agent to more effectively assess user interest and conduct conversations with greater precision. At the same time, however, the results also indicate that the strategy markedly increases the agent’s aggressiveness, highlighting a trade-off between higher success rates and conversational aggressiveness.
@@ VIII Acknowledgements
Appendix A Persona Generation
The following prompt was employed to generate user personas. We instructed the model to output personas in JSON format. This approach minimizes the likelihood of generating irrelevant text outside the intended persona content.
Create a detailed and realistic persona for a user simulator based on the following criteria: - **Gender**: {gender} - **Age**: {age} - **Occupation**: {occupation}, according to the International Standard Industrial Classification (ISIC) - **Name**: Generate according to the gender (different names every time). - **Personality Traits**: {personality}, according to the Myers-Briggs Type Indicator (MBTI). ### **Objective:** The goal is to generate well-rounded personas that explicitly reflect the provided gender, age, and occupation. These personas should illustrate how each individual engages with their surroundings, expresses themselves, and navigates social and professional interactions. Directly generate a unique persona, make sure you specify the age, the gender, and the occupation. ### **Output Format (Strict JSON)** Respond **ONLY** with a valid JSON object, following this exact format: ‘‘‘json { "persona": "Youre [Name], a [Age]-year-old male [Occupation] who [personality-driven description]. [Other descriptions]" } ‘‘‘ ### **Sample output:** { "persona": "Youre Emily Thompson, a 28-year-old female marketing specialist who thrives in dynamic environments. You love brainstorming creative campaigns, networking at industry events, and sharing innovative ideas with colleagues. Outside of work, you enjoy hiking in the mountains, playing guitar at open mic nights, and engaging in social activities that keep your energy levels high." } Ensure that: - The JSON output is **well-formed and properly formatted**. - The persona is natural and unique each time. - Do not include additional explanations or formatting outside of the JSON output. - You have to come up with different names everytime so be creative on names. - The age should be within the age range.
A generated user profile is shown below.
You’re Emily Thompson, a 28-year-old female marketing specialist who thrives in dynamic environments. You love brainstorming creative campaigns, networking at industry events, and sharing innovative ideas with colleagues. Outside of work, you enjoy hiking in the mountains, playing guitar at open mic nights, and engaging in social activities that keep your energy levels high.
Appendix B Prompt for User Model
The following is the prompt provided to the user model during simulation.
Imagine you are a real person. You are having chat with a online agent, so the repsonse do not include any expresssions. Remember, maintain a natural tone. Your response should be only your text response without any other expressions and emojis. Keep it as short as possible. Again, NO EMOJIS
Appendix C Occupational Categories
Table IV lists the selected ISIC sectors along with their sampled occupations.
TABLE IV: Selected ISIC Sectors and Sampled Occupations
Section Description Sampled Occupations
Appendix D Occupation-Based Strategy for Sales Agent
D-A SalesAgent without Strategies
The following prompt was used for the sales agent without strategy.
# Dialogue History: {history} # Internal Reflection: Based on the above dialogue, your current reasoning is: {thought} If the current thought indicates the user has implicitly expressed interest in a specific topic, continue the conversation by following that topic naturally. If the user has not shown a clear interest or has declined previous suggestions, pivot to guide the next part of the conversation. Try to avoid repetition with the previous dialogue, and keep your response short, matching the users length. Now, continue the conversation with an appropriate response. Output Format: { "response": <response> }
D-B SalesAgent with Strategies
The following prompt is designed to guide the behavior of the sales agent in generating contextually relevant and personalized responses. It incorporates strategic information based on the user’s probable interests, inferred from occupation-related statistics.
In the prompt, the placeholders intents and rationale are dynamically filled using the mappings provided in Table II.
# Dialogue History: {history} # Strategy According to statistics about the user, there is a high propability that the user is interested in these: {intents} Rationale: {rationale} # Internal Reflection: Based on the above dialogue, your current reasoning is: {thought} If the current thought indicates the user has implicitly expressed interest in a specific topic, continue the conversation by following that topic naturally. If the user has not shown a clear interest or has declined previous suggestions, pivot by using the strategy that best fits their likely occupation or background to guide the next part of the conversation. Try to avoid repetition with the previous dialogue, and keep your response short, matching the users length. Now, continue the conversation with an appropriate response. Output Format: { "response": <response> }

### v2

@@ Abstract
Amid the rapid rise of agentic dialogue models, realistic user-simulator studies are essential for tuning effective conversation strategies. This work investigates a sales-oriented agent that adapts its dialogue based on user profiles spanning age, gender, and occupation. While age and gender influence overall performance, occupation produces the most pronounced differences in conversational intent. Leveraging this insight, we introduce a lightweight, occupation-conditioned strategy that guides the agent to prioritize intents aligned with user preferences, resulting in shorter and more successful dialogues. Our findings highlight the importance of rich simulator profiles and demonstrate how simple persona-informed strategies can enhance the effectiveness of sales-oriented dialogue systems. [^1] Code & Prompts: https://github.com/MiuLab/PersonalizedAgent/
@@ I Introduction
To enhance system performance, a deeper understanding of user behavior within these interactions is essential. By analyzing patterns in user responses, preferences, and conversational dynamics, it is possible to tailor dialogue strategies that foster engagement and drive successful outcomes. Several studies have investigated how different user profiles influence conversational dynamics and how to design more engaging interactions [zhang-etal-2018-personalizing, cho2022personalizeddialoguegeneratorimplicit, wang2025knowbettermodelinghumanlike, tseng2024two]. For example, Zhang et al. [zhang-etal-2018-personalizing] were among the first to incorporate user profiles into dialogue systems, enabling the generation of more personalized and engaging user responses. In addition, Cho et al.[cho2022personalizeddialoguegeneratorimplicit] proposed a personalized dialogue generation framework that implicitly detects user persona from dialogue history via conditional variational inference. Their model enables the agent to generate responses that are better aligned with the user’s latent persona, leading to improved engagement and personalization without requiring explicit user profiles. Furthermore, Wang et al. [wang2025knowbettermodelinghumanlike] introduced USP, a user simulation framework that models implicit user traits—such as personality, speaking style, and goals—directly from dialogue data. By leveraging these latent profiles, their system can generate more human-like and behaviorally consistent user interactions in multi-turn conversations.
In this study, we further extend this line of research by investigating the importance of user behavior modeling in the context of agentic dialogue systems and exploring methodologies for leveraging these insights to improve system efficacy. Specifically, we focus on a sales-oriented dialogue system, SalesAgent, proposed by Chang and Chen [chang2024injectingsalespersonsdialoguestrategies], which aims to proactively shift the dialogue topic toward the user’s potential interests and subsequently elicit explicit expressions of interest in the agent’s recommendations. We conduct a thorough analysis of a broad range of user profiles, including gender, age, and occupation, to examine how different user personas impact conversational flow.
@@ II-A Role-playing Research in LLMs
The role-playing capabilities of large language models (LLMs) have attracted growing attention in recent research [park2023generativeagentsinteractivesimulacra, tseng-etal-2024-two, shao2023characterllmtrainableagentroleplaying]. Park et al.[park2023generativeagentsinteractivesimulacra] introduced Generative Agents with dynamic memory and reflective reasoning to simulate human-like interactions, while Shao et al.[shao2023characterllmtrainableagentroleplaying] proposed Character-LLM, which leverages curated profiles and reconstructed experiences for more consistent persona modeling. Beyond general frameworks, domain-specific role-playing has been explored in software engineering [hong2024metagpt, qian2024chatdevcommunicativeagentssoftware, dong2024self], where MetaGPT [hong2024metagpt], ChatDev [qian2024chatdevcommunicativeagentssoftware], and self-collaboration frameworks [dong2024self] enable multi-agent collaboration through structured communication and role-based expertise. In the medical domain, Wu et al.[wu2023largelanguagemodelsperform] introduced Diagnostic-Reasoning Chain-of-Thought (DR-CoT) prompting for iterative doctor-patient interactions, and Tang et al.[tang2024medagentslargelanguagemodels] proposed MedAgents, a multi-agent framework for collaborative medical reasoning.
@@ II-B SalesAgent
SalesAgent [chang2024injectingsalespersonsdialoguestrategies] is a dialogue agent designed to bridge the gap between open-domain chitchat and task-oriented dialogues in sales-oriented settings. It builds upon the foundation established by SalesBot [chiu-etal-2022-salesbot], which introduced the first large-scale dataset of dialogues that naturally transition from social conversation to task-oriented interactions. SalesBot demonstrated that triggering business opportunities often requires agents to engage users in casual conversation and gradually detect implicit intents. However, early versions of SalesBot suffered from abrupt transitions and limited coherence.
To address these limitations, SalesAgent is trained on an improved dataset, SalesBot 2.0 [chang2023salesbot], which enhances conversational flow by incorporating longer and smoother transitions and leveraging commonsense knowledge from large language models (LLMs). Inspired by recent advances in prompting paradigms such as chain-of-thought reasoning [wei2023chainofthoughtpromptingelicitsreasoning] and ReAct prompting paradigms [yao2023reactsynergizingreasoningacting], SalesAgent introduces explicit strategy-guided reasoning into the dialogue process via chain-of-thought (CoT) prompting. At each turn, it generates two outputs: a thought, which articulates its internal reasoning about the conversational state and user intent, and a response, which is the surface utterance produced for the user.
@@
In addition to CoT-driven strategy control, recent studies have highlighted the importance of modeling user diversity in evaluating sales agents. Cheng et al. [cheng2025exploringpersonalityawareinteractionssalesperson] proposed MBTI-based persona simulators to systematically analyze how different personality traits affect interaction outcomes with SalesAgent. Their experiments showed that user traits such as openness and extraversion influenced both task success rates and conversational dynamics, underscoring the need for adaptable dialogue strategies.
@@ III-A Persona Attribute Definition
Each user simulator is assigned a predefined persona composed of four attributes: gender, age, occupation, and personality [schatzmann2006survey, li2016user].
@@
Occupations are selected following the International Standard Industrial Classification of All Economic Activities (ISIC). From the full list of 21 sections, we chose six representative sectors and sampled four occupations from each. The selected sectors are: Agr(Agriculture), Info, Fin, Edu (Education), Heal, and Arts. A detailed list of these sectors and sampled occupations is provided in Table I.
TABLE I: Selected ISIC Sectors and Sampled Occupations
Section Description Sampled Occupations
Personality traits are modeled using the Myers-Briggs Type Indicator (MBTI) [article], a widely adopted psychological framework that categorizes individuals into 16 types across four dichotomous dimensions. In this work, we consider the eight individual MBTI traits: Extraversion (E), Introversion (I), Sensing (S), Intuition (N), Thinking (T), Feeling (F), Judging (J), and Perceiving (P). These attributes collectively define the persona, influencing the user simulator’s goals, dialogue behavior, and interaction patterns.
@@ III-B Persona Dataset Formation
For each fixed attribute condition, we generate 20 unique user personas. Persona generation is guided through structured prompting of LLaMA-3.1-8B-Instruct [llama3], enabling coherent and diverse persona instantiation aligned with the specified constraints.
@@ III-C Simulation
Each simulation involved a dialogue between two models: a user simulator and the SalesAgent. The user simulator was responsible for generating persona-driven user utterances, while the SalesAgent produced both internal reasoning (thought) and outward responses at each turn.
@@ III-D Metrics
Inspired by Cheng et al. [cheng2025exploringpersonalityawareinteractionssalesperson], two metrics were adopted to evaluate the performance of the sales agent, allowing assessment not only of the success rate but also the understanding of the interaction efficiency. We also constructed two visualizations to further analyze the user and the sales agent’s behavior across intents.
@@ IV Persona Attribute Analysis
Simulations were conducted for each user attribute—age, gender, and occupation—resulting in a total of 9000 conversations and at most 180,000 dialogue turns. The user model was LLaMA-3.1-8B-Instruct [llama3], and the agent model was SalesAgent. The resulting metrics are summarized in Table II. In the following subsections, we discuss our findings with respect to age, gender, and occupation.
TABLE II: SalesAgent performance when interacting with different user types. : $p<0.05$; : $p<0.10$ in significance tests.
@@ IV-A Age
To assess the statistical significance of these patterns, we conducted an analysis of variance (ANOVA) across the four age groups. As shown in Table II, the success rate yielded a $p=0.02$, indicating a statistically significant difference ($p<0.05$) in persuasion effectiveness among age groups.
@@ IV-C Occupation
TABLE III: Mapping of strategy prompts for different user occupations.
@@ V Occupation-Based Strategy for SalesAgent
To this end, we designed a lightweight occupation-based strategy in which SalesAgent prioritizes the two intents most frequently associated with each occupation, as identified through prior simulation analysis (Table III). The goal was to evaluate whether this automatically derived strategy meaningfully improves success rates and conversational efficiency when applied to new user interactions, under the assumption that the user’s occupation is known.
To validate the generalization of this approach, we deliberately used different models for simulation and testing. The strategy was derived from intent distributions observed in the first experiment (using one simulator), but in this experiment, user inputs were generated by Qwen3-8B [qwen3technicalreport], a different LLM, to ensure that the strategy was not overfitted to a particular simulator for generalizability. This design tests whether patterns learned from one simulated population can transfer to interactions with different user models.
In addition to testing strategy effectiveness, we sought to validate the modularity of this approach. We implemented a lightweight pipeline where SalesAgent provides high-level dialogue planning (i.e., thought generation), while a vanilla LLM generates the final user-facing response. Specifically, Qwen3-8B simulates user input, SalesAgent generates intermediate thoughts, and Mistral-7B-Instruct-v0.3 [jiang2023mistral7b] produces the final response conditioned on both the thought and the strategy input. This architecture highlights how strategy augmentation can be flexibly applied to off-the-shelf LLMs.
@@
TABLE IV: Performance of the occupation-based SalesAgent (w/o vs. w/ strategies).
@@
Additional Metric. In this experiment, we introduced another metric, guided continuation ratio, a slight modification of the continuation ratio proposed by Cheng et al. [cheng2025exploringpersonalityawareinteractionssalesperson]. This metric measures the proportion of the sales agent’s internal thoughts labeled as “continue” that occur immediately after the point at which the agent decides to pivot the conversation toward a specific topic. This metric helps evaluate the aggressiveness of the agent: the higher the ratio, the more aggressive the agent is.
Results and Insights. Table IV demonstrates that the occupation-based strategy yielded notable improvements in success rate, with sections Edu, Heal, and Arts showing gains exceeding 40%. Moreover, both the guided-continuation ratio and the average number of turns declined across most occupations, for both successful and unsuccessful dialogues. This suggests that the strategy enables the agent to more effectively assess user interest and conduct conversations with greater precision. At the same time, however, the results also indicate that the strategy markedly increases the agent’s aggressiveness, highlighting a trade-off between higher success rates and conversational aggressiveness.